OpenStudy (anonymous):

help! Simplify (4xy)(2x2y)(3xy)3. 24x2y3 24x6y5 216x3y3 216x6y5

OpenStudy (anonymous):

I think its A..

OpenStudy (anonymous):

whats this? under what math is this ?

OpenStudy (anonymous):

algebra

OpenStudy (anonymous):

(23•32x4y3) i got that :O

OpenStudy (anonymous):

its not A

OpenStudy (anonymous):

Step 1 : Multiply 4xy by 2x2y Multiplying exponential expressions : 1.1 x1 multiplied by x2 = x(1 + 2) = x3 Multiplying exponential expressions : 1.2 y1 multiplied by y1 = y(1 + 1) = y2 Multiplying exponents : 1.3 22 multiplied by 21 = 2(2 + 1) = 23 Equation at the end of step 1 : (23x3y2 • 3xy) • 3 Step 2 : Multiply 23x3y2 by 3xy Multiplying exponential expressions : 2.1 x3 multiplied by x1 = x(3 + 1) = x4 Multiplying exponential expressions : 2.2 y2 multiplied by y1 = y(2 + 1) = y3 Equation at the end of step 2 : (23•3x4y3) • 3 Step 3 : Multiply (23•3x4y3) by 3 Multiplying exponents : 3.1 31 multiplied by 31 = 3(1 + 1) = 32 Final result : (23•32x4y3)

OpenStudy (anonymous):

oh..uh its not on there

OpenStudy (anonymous):

hmm.. oh wow.. at least im trying

OpenStudy (anonymous):

damn!!! Im out

OpenStudy (anonymous):

great it doesnt even have the answer.....well thanks anyway!